OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -2x-12/-2 A. x − 12 B. x + 6 C. −2x + 6 D. x − 6

OpenStudy (zzr0ck3r):

(-2x-12)/-2 = (-2/-2)x + (-12/-2) = x+6